rishavraj (rishavraj):

Integration of (cos^3 x +cos^5 x)/(sin^2 x + sin^4 x ) dx

OpenStudy (anonymous):

\[\int\limits \frac{ \cos^3 x + \cos^5 x }{ \sin^2 x + \sin^4 x } dx\] ?

rishavraj (rishavraj):

yeah

OpenStudy (anonymous):

wait, i'll try to answer this...

rishavraj (rishavraj):

take ur time.. I m too trying to solve that

OpenStudy (anonymous):

it's not working in my software... i'm using microsoft mathematics...

rishavraj (rishavraj):

microsoft mathematics??

rishavraj (rishavraj):

no need to bother about the question.. I was able to solve it..

OpenStudy (anonymous):

Oh, ok. What was your method? I've been trying partial fractions and a complex substitution, but didn't get far with it.

rishavraj (rishavraj):

u will have to do few manipulations... first of all in numerator take cos x common and in denominator take sin x common... put sin x = t so cos x dx = dt